OpenStudy (anonymous):

a wire 77 inches long is cut into two pieces whose lengths have a ratio of 5 to 2. What is the length of each piece? can you show step by step please

OpenStudy (anonymous):

the wire is cut into 2 pieces, say x and y. x and y are related. x+y=77. Also, x/y=5/2. Solve these 2 equations

OpenStudy (anonymous):

22 and 55

OpenStudy (anonymous):

how do u solve the first equation you put?